### v1

@@ 2.2 Moment-Based Reduced-Order Modeling
(8) $\dot{r}(t)$ $=\bar{f}(r(t),u(t))$
$y_{r}(t)$ $=\bar{h}(r(t))$
with $r(t)\in\mathbb{R}^{v}$, with $v<<n$, $u(t)\in\mathbb{R}^{m}$, $\bar{f}$ is a $C^{r}$ mapping $\bar{f}:\mathbb{R}^{v}\times\mathbb{R}^{m}\to\mathbb{R}^{v}$ satisfying $\bar{f}(0,0)=0$. The output $y_{r}(t)\in\mathbb{R}^{p}$ is measured by the $C^{r}$ mapping $\bar{h}:\mathbb{R}^{v}\to\mathbb{R}^{p}$.
Reduced-order modeling by moment matching involves constructing a system where the observed output $y_{r}(t)$ is in some sense similar to the output of the system $y(t)$ when the input to each system is given by the signal generator. In particular, we are interested in constructing systems where the dimension of the reduced-order model is much smaller, $v<<n$. The main objective behind this is to create smaller models that preserve the behaviour of the system when subjected to signals of interest.
We will use the following definition of a reduced-order model based on [3]:
@@
The system described by (8) is a model at $s(\omega)$ of the system (1) if it has the same moment at $s(\omega)$ as (1). In this case, it is said to match the moment of the system at $s(\omega)$. Furthermore, system (8) is a reduced-order model of system (1) if $v<n$.
@@
(9) $\frac{\partial p}{\partial\omega}s(\omega)$ $=\bar{f}(p(\omega),\ell(\omega))$
@@
Under these assumptions, the steady-state output of (8) is given by
(10) $\lim_{t\to\infty}|y_{r}(t)-\bar{h}(p(\omega(t)))|=0.$
Therefore, the reduced-order model (8) matches the moments of the (1) at $s(\omega)$ if the mapping that solves (9) also satisfies
@@
Figure 1 depicts how moment matching corresponds to the matching of the steady-state responses of the full-order model and reduced-order model.
@@
[…]
@@
Figure 1: Diagrammatic illustration of the concepts of moment matching in reduced-order modeling. The signal generator generates a control input to the higher order model and the reduced-order model, and moment-matching corresponds to the exact matching of their steady-state responses.
@@
In particular, the choice of $p(\omega)=\omega$ gives us a family of reduced-order models
@@
where the mapping $\bar{g}$ is a free parameter, which can be chosen to ensure the local asymptotic stability about the origin of the autonomous system having vector field
(14) $s(r)-\bar{g}(r,\ell(r)).$
@@
It is convenient to have $\bar{g}$ take the form $\bar{g}(r,\ell(r))=g(r)\ell(r)$ for some $C^{r}$ mapping $g:\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$. In the general case, one suitable mapping can be found by computing
@@
and choosing a mapping $g(r)=G$ for some real-valued matrix $G$ satisfying that $S-GL$ has only eigenvalues with negative real part, which can be accomplished if $(S,L)$ is detectable. The linearization of (14) then simply becomes $S-GL$, and the assumption on the eigenvalues of $S-GL$ ensures that the system is therefore locally exponentially stable.
@@ 3 Approximate solutions via Galerkin expansion
(16) $\frac{\partial\pi}{\partial\omega}s(\omega)$ $=f(\pi(\omega),\ell(\omega)),$
@@
This is in general, a difficult problem to solve, as we have to solve a potentially large-scale, high-dimensional system of coupled nonlinear PDEs. A method of approximating the solution to the PDE system (16) is proposed in this section.
@@ 3.1 Polynomial expansion of the invariant mapping
[…]
(17) $\pi^{N}_{i}(\omega)=\sum_{k=1}^{N}c_{i,k}\phi_{k}(\omega)=\Phi^{N}\boldsymbol{c}_{i}$
where $\Phi^{N}=\left(\phi_{1},\phi_{2},\dots,\phi_{N}\right)$, and where $\{\phi_{k}\}_{k=1,\dots,N}$ satisfy $\phi_{k}\in C^{\infty}(\Omega)$ and belong to a complete family of basis functions in $L^{2}(\Omega)$. For this problem, we are interested in globally supported ansatz functions, such as monomials or orthogonal polynomials. The Galerkin residual method finds a weak solution to the invariant manifold $\pi$ by projecting the mapping onto the space of functions $\{\phi_{k}\}$ and ensuring that the approximation error of the equation is orthogonal to every $\phi_{j}$ in the basis. The method of weighted residuals presented here, and the resulting residual equations shown later, are similar to those presented in [14] and [23].
The coefficients of $c_{i}$ in (17) for $i=1,\dots n$ are found by setting the projection of the error in (16) to zero for all $x\in\Omega$ and imposing the Galerkin residual equation
(18) $\left\langle\frac{\partial\pi_{i}}{\partial\omega}s-f_{i}(\pi(\omega),\ell(\omega)),\phi_{j}\right\rangle_{\Omega}=0,$ $\forall\phi_{j}\in\Phi^{N}$
@@
[…]
@@
(23) $F_{i}(\boldsymbol{c})$ $=\begin{bmatrix}\left\langle\frac{\partial\pi_{i}^{N}}{\partial\omega}s(\omega)-f_{i}(\pi^{N}(\omega),\ell(\omega)),\phi_{1}\right\rangle_{\Omega}\\ \vdots\\ \left\langle\frac{\partial\pi_{i}^{N}}{\partial\omega}s(\omega)-f_{i}(\pi^{N}(\omega),\ell(\omega)),\phi_{N}\right\rangle_{\Omega}\end{bmatrix}.$
@@
(24) $F_{i}(\boldsymbol{c})=\left\langle\frac{\partial\Phi^{N}}{\partial\omega}\cdot s,\Phi^{N}\right\rangle_{\Omega}\boldsymbol{c}_{i}-\left\langle f_{i}\left(\left[\Phi^{N}\boldsymbol{c}_{1},\dots,\Phi^{N}\boldsymbol{c}_{n}\right]^{\top},\ell(\omega)\right),\Phi^{N}\right\rangle_{\Omega}.$
The linear term can be written as a matrix-vector product $\boldsymbol{A}\boldsymbol{c}$, where $A$ is defined as
(25) $\boldsymbol{A}$ $=\left\langle\frac{\partial\Phi^{N}}{\partial\omega}\cdot s,\Phi^{N}\right\rangle_{\Omega}.$
@@
(26) $G_{i}(\boldsymbol{c})=-\left\langle f_{i}\left(\left[\Phi^{N}\boldsymbol{c}_{1},\dots,\Phi^{N}\boldsymbol{c}_{n}\right]^{\top},\ell(\omega)\right),\Phi^{N}\right\rangle_{\Omega}.$
@@
If the function $f$ determining the system’s dynamics is polynomial and sufficiently low-dimensional, for example if it is linear, quadratic, or cubic in $\pi$, the functions $G_{i}$ can be represented with matrices and tensors similar to $\boldsymbol{A}$. In problems where this is the case, it will be useful to construct a matrix or tensor representation of the function in order to accelerate both assembly and iterative computations.
In order to find the coefficients $\boldsymbol{c}$ satisfying $F(\boldsymbol{c})=0$, we employ a Newton iteration, which iterates until $|F(\boldsymbol{c})|_{1}$ is below a given tolerance. To carry out this iteration, we also need to find the Jacobian of the Galerkin residual function. The Jacobian is defined as $JF:\mathbb{R}^{Nn}\to\mathbb{R}^{Nn}\times\mathbb{R}^{Nn}$, with $JF(\boldsymbol{c})=\frac{\partial F}{\partial\boldsymbol{c}}(\boldsymbol{c})$. Based on our formulation in (24), we represent it as
@@
(31) $\frac{\partial G_{i}}{\partial\boldsymbol{c}_{j}}(\boldsymbol{c})=-\left\langle\frac{\partial f_{i}(\pi^{N}(\omega),\ell(\omega))}{\partial\boldsymbol{c}_{j}}\Phi^{N},\Phi^{N}\right\rangle_{\Omega}\bigg{|}_{\pi=\Phi^{N}\boldsymbol{c}}.$
Based on the expression of $F(\boldsymbol{c})$ in (24) and of $JF(\boldsymbol{c})$ in (30), the Newton update reads
(32) $\boldsymbol{c}^{k+1}=\boldsymbol{c}^{k}-JF(\boldsymbol{c}^{k})^{-1}F(\boldsymbol{c}^{k})$
@@
Note that in some cases, depending on the system dynamics under study, the Jacobian $JF$ may become singular. In these scenarios, we replace the inverse Jacobian $JF^{-1}$ with the Moore-Penrose inverse (pseudoinverse). However, since the Jacobian is not full-rank, the convergence of the algorithm to the appropriate approximation $\pi^{N}$ will then depend on having a good initialization for the coefficients $\boldsymbol{c}^{0}$.
@@ 3.2 Polynomial basis
The construction of the polynomial basis for our approximation largely follows the procedure used in [23, Section 4.1]. In order to construct the basis functions to be used in the Galerkin approximation over the $d$-dimensional space $\Omega$, we must first select a maximum degree $M$ and a one-dimensional basis $\varphi_{M}:\mathbb{R}\to\mathbb{R}^{M}$. In our case, we will use the monomial basis
(33) $\varphi(\omega)=\left(\omega,\omega^{2},\dots,\omega^{M}\right).$
@@
In order to generate the multi-dimensional basis, we use the tensor product of the one-dimensional basis and eliminate those terms whose total degree is higher than the prescribed maximum degree $M$. The basis is then given by
(34) $\Phi^{N}=\left\{\phi\in\bigotimes_{j=1}^{d}\varphi(\omega_{j}),\text{ where }deg(\phi)\leq M\right\}.$
The number of basis functions $N$ is related combinatorially to the maximum degree $M$, as
(35) $N=\sum_{m=1}^{M}\binom{d+m-1}{m}=\frac{M+1}{d}\binom{M+d}{M+1}-1.$
@@
(36) $\phi_{i}=\prod_{j=1}^{d}\phi_{i}^{j}(x_{j})=\prod_{j=1}^{d}\omega_{j}^{\nu_{i,j}}\text{ where }\sum_{j=1}^{d}\nu_{i,j}\leq M$
for coefficients $\nu_{i,j}$.

### v2

@@ 2.2 Moment-Based Reduced-Order Modeling
(12) $\dot{r}$ $=\bar{f}(r,u)$
$y_{r}$ $=\bar{h}(r)$
with $r(t)\in\mathbb{R}^{v}$, with $v{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ll}n$, $u(t)\in\mathbb{R}^{m}$, and where $\bar{f}$ is a $C^{2}$ mapping $\bar{f}:\mathbb{R}^{v}\times\mathbb{R}^{m}\to\mathbb{R}^{v}$ satisfying $\bar{f}(0,0)=0$. The output $y_{r}(t)\in\mathbb{R}^{p}$ is measured by the $C^{2}$ mapping $\bar{h}:\mathbb{R}^{v}\to\mathbb{R}^{p}$.
The goal is to construct a system of dimension $v{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ll}n$ whose output $y_{r}(t)$ matches the steady-state behaviour of $y(t)$ under the same signal generator input. We will use the following definition of a reduced-order model based on [3]:
@@
The system described by (12) is a model at $s(\omega)$ of the system (1) if it has the same moment at $s(\omega)$ as (1). In this case, it is said to match the moment of the system at $s(\omega)$. Furthermore, system (12) is a reduced-order model of system (1) if it is a model at $s(\omega)$ of (1) and $v<n$.
@@
(13) $\frac{\partial p}{\partial\omega}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega)}s(\omega)$ $=\bar{f}(p(\omega),\ell(\omega))$
@@
Under these assumptions, the steady-state output of (12) from an initial condition $\omega_{0}\in W^{0}$, with $r_{0}$ sufficiently close to the centre manifold, satisfies
(14) $\lim_{t\to\infty}|y_{r}(t;\omega_{0},r_{0})-\bar{h}(p(\Xi_{t}^{s}(\omega_{0})))|=0.$
Therefore, the reduced-order model (12) matches the moments of (1) at $s(\omega)$ if the mapping that solves (13) also satisfies
@@
Figure 1 depicts how moment matching corresponds to the matching of the steady-state responses of the full-order model and the reduced-order model.
@@
$u=\ell(\omega)$$y_{ss}(t)=h(\pi(\omega(t)))$$y(t)$ $\Big\Updownarrow$ (Moment Matching) $y_{ss}(t)=\bar{h}(p(\omega(t)))$
@@
Figure 1: Diagrammatic illustration of the concepts of moment matching in reduced-order modeling. The signal generator generates a control input to the high-order model and the reduced-order model, and moment-matching corresponds to the exact matching of their steady-state responses.
@@
In particular, it is shown in [4] that the choice of $p(\omega)=\omega$ gives us a family of reduced-order models
@@
where the mapping $\bar{g}$ is a free parameter, which can be chosen to ensure the local exponential stability about the origin of the autonomous system having dynamics given by
(18) $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r\mapsto s(r)-\bar{g}(r,\ell(r))$}.
@@
It is convenient to have $\bar{g}$ take the form $\bar{g}(r,\ell(r))=g(r)\ell(r)$ for some $C^{2}$ mapping $g:\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$. In the general case, one suitable mapping can be found by computing
@@
and choosing a mapping $g(r)=G$ for some real-valued matrix $G$ such that SGLS-GL only has eigenvalues with negative real part, which can be accomplished if $(S,L)$ is detectable. The linearization of (18) then simply becomes SGLS-GL, and the assumption on the eigenvalues of SGLS-GL ensures that the system is locally exponentially stable.
@@ 3 Approximate solutions via Galerkin expansion
(20) $\frac{\partial\pi}{\partial\omega}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega)}s(\omega)$ $=f(\pi(\omega),\ell(\omega)),$
@@
A numerical scheme for approximating the solution to (20) is developed in this section.
@@ 3.1 Polynomial expansion of the invariant mapping
Proposition 2.7 only guarantees the existence of a solution to system (20) in a neighbourhood of the origin, hence we restrict our analysis to a compact domain $\Omega\subset\mathbb{R}^{d}$ containing the origin. We denote the $i$-th row of the $n$-dimensional mapping $\pi$ by $\pi_{i}$, so $\pi(\omega)=\left(\pi_{1}(\omega),\pi_{2}(\omega),\dots,\pi_{n}(\omega)\right)^{\top}$. For each mapping $\pi_{i}:\mathbb{R}^{d}\to\mathbb{R}$, $i{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=}1,\dots,n$, we consider an expansion $\pi^{N}_{i}$ of the form
(21) $\pi^{N}_{i}(\omega)=\sum_{k=1}^{N}c_{i,k}\phi_{k}(\omega)=\Phi^{N}(\omega)\boldsymbol{c}_{i}$
where $\Phi^{N}=\left(\phi_{1},\phi_{2},\dots,\phi_{N}\right)$, and where $\{\phi_{k}\}_{k=1,\dots,N}$ satisfy $\phi_{k}\in C^{\infty}(\Omega)$ and belong to a complete family of basis functions in $H^{1}(\Omega)$. Note that we choose $H^{1}(\Omega)$ to ensure that $\frac{\partial\phi_{i}}{\partial\omega}\in L^{2}(\Omega)$. Here, $\boldsymbol{c_{i}}=(c_{i,1},\dots,c_{i,k},\dots,c_{i,N})^{\top}\in\mathbb{R}^{N}$ denotes the vector of coefficients associated with the basis functions $\phi_{k}$ for the $i$-th component of $\pi$. For this problem, we are interested in globally supported ansatz functions, such as monomials or orthogonal polynomials. The Galerkin residual method finds a weak solution to the invariant manifold $\pi$ by projecting the mapping onto the space of functions $\{\phi_{k}\}$ and ensuring that the approximation error of the equation is orthogonal to every $\phi_{j}$ in the basis. The method of weighted residuals presented here and the resulting residual equations shown later are similar to those presented in [14] and [23].
The components of $\boldsymbol{c}_{i}$ in (21) for $i=1,\dots,n$ are found by setting the projection of the error in (20) to zero for all $\omega\in\Omega$ and imposing the Galerkin residual equation
(22) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left\langle\frac{\partial\pi_{i}}{\partial\omega}s-f_{i}(\pi,\ell),\phi_{j}\right\rangle_{\Omega}=0,}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{for }j=1,\dots,N}$
@@
(23) $\langle f,g\rangle_{\Omega}=\int_{\Omega}f(\omega)g(\omega)d\omega.$
We note that $\frac{\partial\pi_{i}}{\partial\omega}(\cdot)s(\cdot)-f_{i}(\pi(\cdot),\ell(\cdot))\in L^{2}(\Omega)$ because $\pi$ is smooth by the definition of the centre manifold, $s$ and $f$ are assumed to be at least $C^{2}$, and $\Omega$ is compact.
@@
(27) $F_{i}(\boldsymbol{c})$ $=\begin{bmatrix}\left\langle\frac{\partial\pi_{i}^{N}}{\partial\omega}s-f_{i}(\pi^{N},\ell),\phi_{1}\right\rangle_{\Omega}\\ \vdots\\ \left\langle\frac{\partial\pi_{i}^{N}}{\partial\omega}s-f_{i}(\pi^{N},\ell),\phi_{N}\right\rangle_{\Omega}\end{bmatrix}$.
@@
(28) $F_{i}(\boldsymbol{c})=\left\langle\frac{\partial\Phi^{N}}{\partial\omega}s,\Phi^{N}\right\rangle_{\Omega}\boldsymbol{c}_{i}-\left\langle f_{i}\left(\left[\Phi^{N}\boldsymbol{c}_{1},\dots,\Phi^{N}\boldsymbol{c}_{n}\right]^{\top},\ell\right),\Phi^{N}\right\rangle_{\Omega}.$
[…]
(29) $\boldsymbol{A}$ $=\left\langle\frac{\partial\Phi^{N}}{\partial\omega}s,\Phi^{N}\right\rangle_{\Omega}.$
@@
(30) $G_{i}(\boldsymbol{c})=-\left\langle f_{i}\left(\left[\Phi^{N}\boldsymbol{c}_{1},\dots,\Phi^{N}\boldsymbol{c}_{n}\right]^{\top},\ell\right),\Phi^{N}\right\rangle_{\Omega}.$
@@
If the function $f$ determining the system’s dynamics is a polynomial of sufficiently low degree, for example if it is linear, quadratic, or cubic in $\pi$, the functions $G_{i}$ can be represented with matrices and tensors similar to $\boldsymbol{A}$. In problems where this is the case, it will be useful to construct a matrix or tensor representation of the function in order to accelerate both assembly and iterative computations.
In order to find the coefficients $\boldsymbol{c}$ satisfying $F(\boldsymbol{c})=0$, we employ a Newton iteration, which iterates until $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|F(\boldsymbol{c})\|_{1}$} is below a given tolerance. To carry out this iteration, we also need to find the Jacobian of the Galerkin residual function. The Jacobian is defined as $JF:\mathbb{R}^{Nn}\to{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{Nn\times Nn}}$, with $JF(\boldsymbol{c})=\frac{\partial F}{\partial\boldsymbol{c}}(\boldsymbol{c})$. Based on our formulation in (28), we represent it as
@@
(35) $\frac{\partial G_{i}}{\partial\boldsymbol{c}_{j}}(\boldsymbol{c})=-\left\langle\frac{\partial f_{i}(\pi^{N},\ell)}{\partial\boldsymbol{c}_{j}}\Phi^{N},\Phi^{N}\right\rangle_{\Omega}\bigg|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi^{N}=\Phi^{N}\boldsymbol{c}}}.$
Based on the expression of $F(\boldsymbol{c})$ in (28) and of $JF(\boldsymbol{c})$ in (34), we calculate the Newton increment $\Delta\boldsymbol{c}^{k}$ by solving
(36) $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}JF(\boldsymbol{c}^{k})\Delta\boldsymbol{c}^{k}=F(\boldsymbol{c}^{k}),$}
and the Newton update is then
(37) $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{c}^{k+1}=\boldsymbol{c}^{k}-\Delta\boldsymbol{c}^{k}.$}
@@
Note that in some cases, depending on the system dynamics under study, the Jacobian $JF$ may become singular. In these scenarios, we solve for the Newton increment $\Delta\boldsymbol{c}^{k}$ in the least-squares sense.
@@ 3.2 Polynomial basis
The construction of the polynomial basis for our approximation largely follows the procedure used in [23, Section 4.1]. In order to construct the basis functions to be used in the Galerkin approximation over the $d$-dimensional space $\Omega$, we first select a maximum degree $M$ and construct a basis in a single scalar variable $\omega$ up to the degree $M$. We denote this basis by $\varphi_{M}:\mathbb{R}\to\mathbb{R}^{M}$. In our case, we will use the monomial basis
(38) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi_{M}(\omega)}=\left(\omega,\omega^{2},\dots,\omega^{M}\right).$
@@
In order to generate the multi-dimensional basis, we use the tensor product of the basis in each coordinate and eliminate those terms whose total degree is higher than the prescribed maximum degree $M$. The basis is then given by
(39) $\Phi^{N}=\left\{\phi\in\bigotimes_{j=1}^{d}\varphi_{M}(\omega_{j}),\text{ where }{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\deg}(\phi)\leq M\right\}.$
The number of basis functions $N$ is combinatorially related to the maximum degree $M$. A stars-and-bars argument is used to count the number of monomials of a given total degree, and the summation over all degrees up to $M$ can be evaluated in closed form using the hockey-stick identity, yielding
(40) $N=\sum_{m=1}^{M}\binom{d+m-1}{m}=\binom{M+d}{M}-1.$
@@
(41) $\phi_{i}(\omega)=\prod_{j=1}^{d}\phi_{i}^{j}(\omega_{j})=\prod_{j=1}^{d}\omega_{j}^{\nu_{i,j}}\,\text{ where }\sum_{j=1}^{d}\nu_{i,j}\leq M$
for exponents $\nu_{i,j}$.